1- Suppose there are only 38 students in advance statistics class and final scores of these 38 students are: 10 34 45 55 63 75 class lit 23 34 49 56 64 82 24 36 50 58 65 82 26 41 51 59 71 95- 32 41 52 60 73 33 45 54 62 74 s lin 46 54 63 74 a- Construct grouped frequency distribution table for this dataset with relative and cumulative frequencies (with closed intervals). b- Find mean for this grouped dataset. c- Find median and mode for this grouped dataset. d- Find standard deviation. e- Draw a histogram for this grouped dataset. f- Draw ogive for the grouped dataset. What does cumulative frequency in 3rd d.umalative frequeney in Jrd class indicate? g-According to Chebyshevs Theorem, in which range (or interval) does % of the data values change in?
Answer #1
(a)
CI f Relative f Cumulative f
10 - 30 4 0.105 4
30 - 50 11 0.289 15
50 - 70 15 0.395 30
70 - 90 7 0.184 37
90 - 110 1 0.026 38
Sums = 38 1
(b)
x f fx
20 4 80
40 11 440
60 15 900
80 7 560
100 1 100
Sums = 38 2080
Mean = ∑fx/∑f = 2080/38 = 54.74
(c) Median class is 50 - 70
L = 50, N/2 = 19, CF = 15, F = 15, H = 20
Median = L + ((N/2) - CF) * H/F = 50 + (19 - 15) * 20/15 = 55.33
Modal class is 50 - 70
L = 50, F1 = 15, F0 = 11, F2 = 7, H = 20
Mode = L + ((F1 - F0)/(2F1 - F0 - F2)) * H = 50 + ((15 - 11)/(2 * 15 - 11 - 7)) * 20 = 60
(d)
x f d = x - Mean d^2 fd^2
20 4 -34.74 1206.868 4827.470
40 11 -14.74 217.268 2389.944
60 15 5.26 27.668 415.014
80 7 25.26 638.068 4466.473
100 1 45.26 2048.468 2048.468
Sums = 38 14147.37
Standard deviation = √[∑fd^2 / ∑f] = √ [14147.37/38] = 19.295

(e) Histogram

14 12 10 70-90 90 110 it 50-70 ญา 10-30 30-50 Cl

(f)
Less than CF
30 4
50 15
70 30
90 37
110 38

5 4 3 050505050 4332211

Cumulative frequency in the third class indicates the number of students who scored less than 70
(g) 1 - (1/k)^2 = 0.75
Solving, k = 2
Mean - 2 * Standard deviation = 54.74 - 2 * 19.295 = 16.15
Mean + 2 * Standard deviation = 54.74 + 2 * 19.295 = 93.33
The interval is [16.15, 93.33]